OpenStudy (ray10):

Caluclate the change in volume of an aluminium sphere of radius 10cm, if it is heated from \[\large 0^{o}\] to \[\large 60^{o}C\] \[\large \alpha_{aluminium} = 23 \times 10^{-6} \]

OpenStudy (ray10):

@AllTehMaffs

OpenStudy (anonymous):

what's alpha and what are the units on it? ^^

OpenStudy (anonymous):

linear expansion coefficient?

OpenStudy (ray10):

so the formula to work this out is: \[\large \Delta V = 3\alpha \times V_{0} \times \Delta\] where Alpha is the expansion properties of the desired material, (aluminium)

OpenStudy (ray10):

from what I can see in my notes, it is unit less, so a coefficient yes

OpenStudy (anonymous):

it should be in units of \[K^{-1} \] there's a temperature in there too I hope? \[ 3 \ \alpha \ V_0 \Delta T=\Delta V\]

OpenStudy (ray10):

oh gosh I forgot to add the T when I was showing you it, yes there is meant to be a T there, sorry

OpenStudy (anonymous):

Then plug and chug!!! \[ V_0 = V_{sphere} = 4\pi r^3/3\] \[ \Delta T = 60ºC - 0º = 60 K\] \[\alpha = 26 x 10^{-6} K^{-1} \]

OpenStudy (anonymous):

No worries - figured that was what was attached to that lonesome delta ^^

OpenStudy (ray10):

I can get to that part thank you :) it's the final answer I get stuck on :(

OpenStudy (anonymous):

That is the final answer.... :/

OpenStudy (anonymous):

you're solving for \[ \Delta V\]

OpenStudy (ray10):

I get \[\large 3 \times (23 \times 0.01^{-6}) \times (\frac{4}{3}\pi10^{3}) \times 60\] = 0.00000001734

OpenStudy (ray10):

well I use 0.01 as \[\large \frac {4}{3}\pi(0.01)^{3})\]

OpenStudy (anonymous):

What's the answer supposed to be?

OpenStudy (ray10):

according to my answer sheet, it must be \[\large 17.3cm^{3}\]

OpenStudy (anonymous):

Then it should be 10 for the radius

OpenStudy (ray10):

so will the radius be in centimeters then?

OpenStudy (anonymous):

yeah

OpenStudy (ray10):

so no need to convert to 0.01?

OpenStudy (anonymous):

nope nope - that's why i's so tiny. Cubic things get tiny quick at different orders of magnitude

OpenStudy (ray10):

oh I see no :D so in these cases, I shall always be working with centimeters ? :)

OpenStudy (anonymous):

I'm not really sure for every case - this one if it's in meters then it would be quite tiny. If you wrote the answer down like you had it before \[1.734 x 10^{-8}m^3\] you'd still be right. That's an equivalent answer - just make sure to keep track of what you're using.

OpenStudy (ray10):

ah now I see :) Thank you for your help @AllTehMaffs !! I'll keep a more watchful eye on these problems and look at them the way you mentioned to me :)

OpenStudy (anonymous):

You're welcome ^_^ Physics is all just one big game of "keep track of the units!". If you're ever in a bind, ignore numbers first and see what the raw dimensional analysis of the problem yields. ^_^